OpenStudy (anonymous):

Help please :c Being a smart financial planner, you want to figure out how many months it will be until your principal is paid down to $10,000.00. Solve for t and show all of your work. Note that t will be negative because the number of months will decrease the principal. I started with 10,000=39145(1+.03/12)^12t I ended up with in(10000/39145)/12in(1.0025) =t My only problem is I keep getting t=0.021235284081 for an answer which I believe is wrong.

OpenStudy (anonymous):

@ganeshie8 Any ideas?

OpenStudy (anonymous):

@phi Help?

OpenStudy (anonymous):

Did I get the function wrong when I was isolating t?

OpenStudy (phi):

you should get \[ \frac{10000}{39145} = 1.0025^{12 t } \] \[ \ln\left( \frac{10000}{39145} \right) = 12 \ln(1.0025) t \] \[ \frac{ \ln\left( \frac{10000}{39145} \right)}{12 \ln(1.0025)} = t \]

OpenStudy (phi):

ln means "natural log" . You could also use log (base 10) you will get a negative number.

OpenStudy (anonymous):

That exactly what I have then I solved it using Mathway and t=0.021235284081 is what I get, I'm just no so confident in myself that it's right.

OpenStudy (anonymous):

So I'd re-do the problem but change where "in" is to "log(10)"?

OpenStudy (phi):

You should get a negative number, not 0.0212 do you have a calculator? or use google: log(10000/39145)= to get the top number

OpenStudy (anonymous):

okay let me solve it one sec

OpenStudy (anonymous):

I got -0.5926762974

OpenStudy (phi):

now find 12*log(1.0025)

OpenStudy (anonymous):

I got 0.0130125755

OpenStudy (anonymous):

So would I divide them both together?

OpenStudy (phi):

yes, now do -0.5926762974/0.0130125755

OpenStudy (phi):

type into google -0.5926762974 / 0.0130125755 =

OpenStudy (anonymous):

I got -45.5464252561

OpenStudy (anonymous):

So that's my answer?

OpenStudy (phi):

yes, -45.55 months, or -46 months rounded

OpenStudy (anonymous):

Thank you so much man!!

OpenStudy (phi):

yw